OpenStudy (anonymous):

If sin y° = 9/c. and tan y° = 9/d. ,what is the value of cos y°?

OpenStudy (welshfella):

use the identity tan y = sin y / cos y so plug in the given values for the sin and the tan and fond the cos

OpenStudy (welshfella):

tan y = sin y / cos y cross multiply tan y cos y = sin y divide both sides by tan y;- cos y = sin y / tan y now plug in the given values to find cos y

OpenStudy (michele_laino):

Please try to apply this identity: \[\cos x = \frac{{\sin x}}{{\tan x}}\]